# When Can Model-Free Reinforcement Learning be Enough for Thinking?

### v1

@@ 4 Policy Initialization Determines Emergence of Thinking
The inequality $v_{\pi}(s,\tau^{\prime})>v_{\pi}(s,\tau)$ immediately follows from Theorem 5 The inequality $v_{\pi}(s,\tau^{\prime\prime})>v_{\pi}(s,\tau^{\prime})$ follows from the same logic as the proof of Theorem 5 except applied to improving the policy in $(s,\tau^{\prime})$. ∎
If each step of thinking in $s$ improves the policy, $\pi(s,\cdot)$, when should thinking terminate? From the proof of Theorem 5, we can see that thinking will terminate when the increase in value from thinking another step no longer compensates for the discounting of value caused by waiting a step to begin taking environmental actions. This observation connects the “when to think" decision directly to the agent’s objective to maximize its total reward.
@@ 5 Language Generation as a Thought MDP
Our work takes inspiration from recent work on large reasoning models (LRMs) that “think" by generating additional text that is not part of the final answer but that somehow serves to improve the final answer. In this section, we review two approaches to imbue LLMs with reasoning capabilities and describe how they can be viewed as instantiating thought MDPs. We then show that forcing the LLM to reason (in a manner similar to zero-shot chain-of-thought (Kojima et al., 2023)) increases the expected return from a given state. Thus, Theorem 5 predicts that model-free RL applied to this thought MDP would lead to thinking, which is in fact what recent work has found.
@@
The main difficulty in mapping language generation to a thought MDP is that thought states are intertwined with environment states in the typical MDP formulation. Similarly, thought actions are simply from the same space as environment actions. To distinguish these components, we redefine the environment state as just the query and tokens that function as part of the query response. Next, we define environment actions as just outputs that will affect how the overall response is evaluated and thought actions as the outputs that will not. Thus, the determination of whether one output is considered an environment or a thought action will depend upon the current context up to that decision. Finally, the thought state at time $t$ consists of the already produced tokens in $y_{1:t}$ that will not affect how the response is evaluated.
Different approaches to inducing reasoning in LLMs use different schemes for determining which tokens are part of the final response and which only serve to improve the final response. For example, Guo et al. (2025) augment the output vocabulary with two special functional tokens, <think> and </think>. In addition to the sparse verifier reward, Guo et al. train DeepSeek-R1 with reward shaping to encourage valid thinking blocks in which <think> is followed by </think>. Output text in thinking blocks is not part of the final response that is passed to a verifier to determine reward and would thus constitute the thought state. Zero-shot prompting is another approach to encourage thinking-like behavior by appending “Let’s think step-by-step" to the query Kojima et al. (2023). In this approach, an answer parser is used to separate the response (environment actions) from the additional prompt and subsequent reasoning tokens (thought actions). The outputs that are discarded by the parser would constitute the thought state.
@@
Recent work from DeepSeek (and possibly OpenAI) has shown that thinking-like behavior can emerge from RL. Based on our theory, we hypothesize that a pre-condition for this result is that thought actions increase $v_{\pi}(s,\tau)$ by changing the thought state $\tau$. Because we lack a value function for each LLM, we instead approximate $v_{\pi}(s,\tau)$ with the Monte Carlo return or, equivalently, the accuracy of the LLM’s response. Our hypothesis is that forcing the LLM to think will increase this value. Note that we use the Monte Carlo return as opposed to just looking at the immediate probability of the correct response (or its perplexity) to allow for the possibility that the LLM will generate its own reasoning and then answer correctly. To test our hypothesis, we take different pre-trained LLMs and apply them to add series of five four-digit numbers. We use models of various sizes from the following model familes: Qwen-2.5 (Yang et al., 2024; Qwen Team, 2024), Tulu-2 (Ivison et al., 2023), LLama-2 (Touvron et al., 2023), Gemma-3 (Gemma Team, 2025), and Mistral (Jiang et al., 2023). We test two conditions with each model: “No Thinking" and “Thinking." Under the “No Thinking" condition, the prompt is “Compute the sum: [a] + [b] + [c] + [d] + [e] = " where [a], [b], [c], [d], and [e] correspond to four-digit integers. Under the “Thinking" condition, we append “[a + b] + [c] + [d] + [e] = [a + b + c] + [d] + [e] = [a + b + c + d] + [e] = " to this prompt, where [a + b] and [a + b + c] denote the partial sums a+b and a+b+c. In essence, we force the model to first think under the “Thinking" condition. We constuct 1000 “Thinking" and 1000 “No Thinking" prompts like these using 1000 different sequences of four 4-digit integers, each generated uniformly at random (see Appendix B for example prompts). Table 1 shows the average accuracy for each model. For all models, we see that appending the thinking tokens increases accuracy, which corresponds to increasing $v_{\pi}(s,\tau)$ by changing $\tau$. While we do not further apply model-free RL to try and learn to think in this way, our theory and these results predict that these models are primed for thinking to further emerge as a strategy.
@@ 6 A Non-Language Thought MDP
Our work takes inspiration from work on LLMs but aspires to understand how RL can lead to thinking in domains beyond language generation. One of the most important open questions is where do thought MDPs and initial thought-conditioned policies come from outside of LLMs? In this section, we hypothesize about generalized ingredients for a setting where thinking emerges as a useful strategy for reward maximization. Specifically, we hypothesize that the key ingredients may be multi-task pre-training coupled with the agent having the ability to manipulate its own internal state to activate pre-trained abilities. LLMs fit this hypothesis: self-supervised pre-training enables LLM agents to generate responses to many different contexts and language actions enable RL-finetuning to manipulate the input context to activate these existing capabilities when doing so will produce reward. To test this hypothesis, we set out to construct a non-language toy domain with these characteristics and to see if thinking could enable the agent to reach a higher return compared to non-thinking agents.
We design a 5x5 gridworld where the agent can move in the cardinal directions or output one of three special actions, ‘A’, ‘B’, or ‘C’ that have no direct effect on the environment. Each special action corresponds to a possible task in the gridworld: ‘A’ corresponds to navigating to the lower right corner, ‘B’ corresponds to the upper left corner, and ‘C’ corresponds to going to the lower right corner and then to the upper left corner. Note that these special actions do not explicitly have these effects in the domain but will be used as thoughts that can trigger these behaviors after pre-training. At the start of each episode, the agent receives one of these special actions as part of its observation.
We pre-train an initial policy through behavior cloning on an agent that “plays" in the gridworld. Specifically, this agent samples a new special action with probability 0.25 and subsequently acts optimally for the corresponding task until it succeeds or samples a new task. During pre-training we only consider special actions ‘A’ and ‘B’ so that the agent has no experience with ‘C’. The result of this procedure is a policy that wanders back and forth between the upper left and lower right corner while randomly changing its mind about where it is going. For the policy, we use a causal transformer (Vaswani et al., 2017) that conditions the probability of its action choice on the full episode history up to time $t$. As actions are included in this history, the special actions can manipulate the internal state of the agent even though they do not affect the environment state.
@@
Figure 2: Learning curves for the four agents we train in the gridworld environment. The vertical axis gives the success rate for first navigating to the bottom right and then the top left corner. The horizontal axis is the iteration of policy improvement (200 episodes are collected at each iteration). We run 5 trials for each learning agent and shading indicates a 95% bootstrap confidence interval.
In Figure 2 we see that Pretrained-Think learns significantly faster than the other agents. The sparse reward makes learning from scratch impossible with or without thinking. Pretrained-NoThink also begins to learn the task, though it takes many more iterations to do so. This result was initially surprising to us and prompted further investigation on why pre-training helped even without thinking. The apparent reason is that sequences of environment actions can also serve to trigger the pre-trained policy’s sub-task behavior in a way similar to thought actions. For example, if the agent moves down 2-3 times in a row then the agent will tend to keep moving toward goal ‘A’ at the bottom right. Nevertheless, it is more difficult to discover these sequences compared to taking a single thought action.
We also investigate if thinking is actually what is learned by Pretrained-Think. Figure 2 shows that the agent initially takes a thought action about 30% of the time and this fraction falls as it learns and then stabilizes at about 15%. Optimal episodes are 14 steps long and the optimal amount of thinking – barring the agent figuring out the true optimal policy – is 2 times which is approximately 15% of all time-steps at convergence. We observe a trained agent and confirm that it first outputs ‘A,’ navigates to the bottom right, outputs ‘B,’ and then navigates to the top left.
@@ 7 Limitations for Future Work
This work studied the question of why thinking actions could be useful to an RL agent even though they leave their environment state unchanged and produce no reward. We formalized this abstract notion of thinking with the thought MDP model, but left open the big question of how to define the thought states, actions, and thought dynamics in other RL problems where thinking may be useful. Language generation is one example, but it is an open question as to how thought MDPs might arise outside the language domain. There is also the question of where existing thought-conditioned policies come from, as our work showed their existence to be a key enabler of emergent thinking. In Section 6, we explored the hypothesis that multi-task pre-training could be a key ingredient but this experiment was only designed as a proof-of-concept. In the future, it would be interesting to extend this exploration to more complex and realistic domains such as robots using pre-trained vision-language-action models.

### v2

@@ 4 Policy Initialization Determines Emergence of Thinking
The inequality $v_{\pi}(s,{\tau}^{\prime})>v_{\pi}(s,\tau)$ immediately follows from Theorem˜5. The inequality $v_{\pi}(s,{\tau}^{\prime\prime})>v_{\pi}(s,{\tau}^{\prime})$ follows from the same logic as the proof of Theorem˜5 except applied to improving the policy in $(s,{\tau}^{\prime})$. ∎
If each step of thinking in $s$ improves the policy, $\pi(s,\cdot)$, when should thinking terminate? From the proof of Theorem˜5, we can see that thinking will terminate when the increase in value from thinking another step no longer compensates for the discounting of value caused by waiting a step to begin taking environmental actions. In this way, thought MDPs directly tie the “when to think" decision to the objective of reward maximization.
@@ 5 Language Generation as a Thought MDP
Our work takes inspiration from recent work on large reasoning models (LRMs) that “think" by generating additional text that is not part of the final answer but that somehow serves to improve the final answer. In this section, we review two approaches to imbue LLMs with reasoning capabilities and describe how they can be viewed as instantiating thought MDPs. We then show that forcing the LLM to reason (in a manner similar to zero-shot chain-of-thought (Kojima et al., 2023)) increases the expected return from a given state. Thus, Theorem˜5 predicts that model-free RL applied to this thought MDP would lead to thinking, which is in fact what recent work has found.
@@
The main difficulty in mapping language generation to a thought MDP is that thought states are intertwined with environment states in the typical MDP formulation for LLMs. Similarly, thought actions are simply from the same space as environment actions. To distinguish these components, we redefine the environment state as just the query and tokens that function as part of the query response. Next, we define environment actions as just outputs that will affect how the overall response is evaluated and thought actions as the outputs that will not. Thus, the determination of whether one output is considered an environment or a thought action will depend upon the current context up to that decision. Finally, the thought state at time $t$ consists of the already produced tokens in $y_{1:t}$ that will not affect how the response is evaluated.
Different approaches to inducing reasoning in LLMs use different schemes for determining which tokens are part of the final response and which only serve to improve the final response. For example, Guo et al. (2025) augment the output vocabulary with two special functional tokens, <think> and </think>. In addition to the sparse verifier reward, Guo et al. train DeepSeek-R1 with reward shaping to encourage valid thinking blocks in which <think> is followed by </think>. Output text in thinking blocks is not part of the final response that is passed to a verifier to determine reward and would thus constitute the thought state. Zero-shot prompting is another approach to encourage thinking-like behavior by appending “Let’s think step-by-step" to the query (Kojima et al., 2023). In this approach, an answer parser is used to separate the response (environment actions) from the additional prompt and subsequent reasoning tokens (thought actions). The outputs that are discarded by the parser would constitute the thought actions.
@@
Recent work from Guo et al. (2025) has shown that thinking-like behavior can emerge from RL. Based on our theory, we hypothesize that a pre-condition for this result is that thought actions increase $v_{\pi}(s,\tau)$ by changing the thought state $\tau$. Because we lack a value function for each LLM, we instead approximate $v_{\pi}(s,\tau)$ with the Monte Carlo return or, equivalently, the accuracy of the LLM’s response. Our hypothesis is that forcing the LLM to think will increase this value. Note that we use the Monte Carlo return as opposed to just looking at the immediate probability of the correct response (or its perplexity) to allow for the possibility that the LLM will generate its own reasoning and then answer correctly. To test our hypothesis, we take different pre-trained LLMs and apply them to add series of five four-digit numbers. We use models of various sizes from the following model familes: Qwen-2.5 (Yang et al., 2024; Qwen Team, 2024), Tulu-2 (Ivison et al., 2023), LLama-2 (Touvron et al., 2023), Gemma-3 (Gemma Team, 2025), and Mistral (Jiang et al., 2023). We test two conditions with each model: “No Thinking" and “Thinking." Under the “No Thinking" condition, the prompt is “Compute the sum: [a] + [b] + [c] + [d] + [e] = " where [a], [b], [c], [d], and [e] correspond to four-digit integers. Under the “Thinking" condition, we append “[a + b] + [c] + [d] + [e] = [a + b + c] + [d] + [e] = [a + b + c + d] + [e] = " to this prompt, where [a + b] and [a + b + c] denote the partial sums a+b and a+b+c. In essence, we force the model to first think under the “Thinking" condition. We constuct 1000 “Thinking" and 1000 “No Thinking" prompts like these using 1000 different sequences of four 4-digit integers, each generated uniformly at random (see Appendix˜D for example prompts). Table˜1 shows the average accuracy for each model. For all models, we see that appending the thinking tokens increases accuracy, which corresponds to increasing $v_{\pi}(s,\tau)$ by changing $\tau$. While we do not further apply model-free RL to try and learn to think in this way, our theory and these results predict that these models are primed for thinking to further emerge as a strategy.
@@ 6 A Non-Language Thought MDP
Our work takes inspiration from work on LLMs but aspires to understand how RL can lead to thinking in domains beyond language generation, possibly including agents that think in sequences of images (Ghazanfari et al., 2025; Wiedemer et al., 2025) or more abstract spaces (Hao et al., 2024). One of the most important open questions is where do thought MDPs and initial thought-conditioned policies come from outside of LLMs? In this section, we hypothesize about generalized ingredients for a setting where thinking emerges as a useful strategy for reward maximization. Specifically, we hypothesize that the key ingredients may be multi-task pre-training coupled with the agent having the ability to manipulate its own internal state to activate pre-trained abilities. LLMs fit this hypothesis: self-supervised pre-training enables LLM agents to generate responses to many different contexts and language actions enable RL-finetuning to manipulate the input context to activate these existing capabilities when doing so will produce reward. To test this hypothesis, we set out to construct a non-language toy domain with these characteristics and to see if thinking could enable the agent to reach a higher return compared to non-thinking agents. Code for these experiments is available at https://github.com/prediction-action-lab/thinking-as-control.
We design a 5x5 gridworld where the agent can move in the cardinal directions or output one of three special actions, ‘A’, ‘B’, or ‘C’ that have no direct effect on the environment. Each special action corresponds to a possible task in the gridworld: ‘A’ corresponds to navigating to the lower right corner, ‘B’ corresponds to the upper left corner, and ‘C’ corresponds to going to the lower right corner and then to the upper left corner. Note that these special actions do not explicitly have these effects in the domain but will become associated with these behaviors through pre-training. [^2] In Appendix E, we also include results for the case where tasks ‘A’ and ‘B’ are slightly different from the sub-tasks needed to complete task ‘C.’ At the start of each episode, the agent receives one of these special actions as part of its observation.
We pre-train an initial policy through behavior cloning on an agent that “plays" in the gridworld. Specifically, at each time-step, this agent samples a new special action with probability 0.25 and subsequently acts optimally for the corresponding task until it succeeds or samples a new task. During pre-training we only consider special actions ‘A’ and ‘B’ so that the agent has no experience with ‘C’. The result of this procedure is a policy that wanders back and forth between the upper left and lower right corner while randomly changing its mind about where it is going. For the policy, we use a causal transformer (Vaswani et al., 2017) that conditions the probability of its action choice on the full episode history up to time $t$. As actions are included in this history, the special actions can manipulate the internal state of the agent even though they do not affect the environment state.
@@
Figure 2: Mean learning curves for the four agents we train in the gridworld environment. The vertical axis gives the success rate for first navigating to the bottom right and then the top left corner. The horizontal axis is the iteration of policy improvement (200 episodes are collected at each iteration). We run 5 trials for each learning agent and shading indicates a 95% bootstrap confidence interval. Light lines show individual training runs.
In Figure˜2 we see that Pretrained-Think learns significantly faster than the other agents. The sparse reward makes learning from scratch impossible with or without thinking. Pretrained-NoThink also begins to learn the task, though it takes many more iterations to do so. This result was initially surprising to us and prompted further investigation on why pre-training helped even without thinking. The apparent reason is that sequences of environment actions can also serve to trigger the pre-trained policy’s sub-task behavior in a way similar to thought actions. For example, if the agent moves down 2-3 times in a row then the agent will tend to continue to goal ‘A’ at the bottom right. Nevertheless, it is more difficult to discover such sequences compared to taking a single thought action.
We also investigate if thinking is actually what is learned by Pretrained-Think. Figure˜2 shows that the agent initially takes a thought action about 30% of the time and this fraction falls as it learns and then stabilizes at about 15%. Optimal episodes are 14 steps long and the optimal amount of thinking – barring the agent figuring out the true optimal policy – is 2 times which is approximately 15% of all time-steps at convergence. We observe a trained agent and confirm that it first outputs ‘A,’ navigates to the bottom right, outputs ‘B,’ and then navigates to the top left.
Finally, we note that thought actions do not go away as the policy converges – in contrast to Proposition˜4 which states that the optimal policy only takes environment actions. We suspect that this finding is due to sampling error that can cause poor convergence of on-policy policy gradient methods learning from finite samples (Corrado and Hanna, 2023). As the agent discovers the strategy to first think and then act, it puts increasingly less probability on immediately taking the optimal action. Consequently, it never accidently discovers the benefit of taking immediate action to reach the goal a step or two earlier.
@@ 7 Limitations for Future Work
This work studied the question of why thinking actions could be useful to an RL agent even though they leave their environment state unchanged and produce no reward. We formalized this abstract notion of thinking with the thought MDP model, but left open the big question of how to define the thought states, actions, and thought dynamics in other RL problems where thinking may be useful. Language generation is one example, but it is an open question as to how thought MDPs might arise outside the language domain. There is also the question of where existing thought-conditioned policies come from, as our work showed their existence to be a key enabler of emergent thinking. In Section˜6, we explored the hypothesis that multi-task pre-training could be a key ingredient but this experiment was only designed as a proof-of-concept. In the future, it would be interesting to extend this exploration to more complex and realistic domains such as robots using pre-trained vision-language-action models or visual reasoning tasks using video models (Wiedemer et al., 2025).
@@ 8 Conclusion
Acknowledgments
We thank Xiaojin Zhu and Adam Labiosa for much discussion and comments that improved the final version of this work. This work took place in the Prediction and Action Lab (PAL) at the University of Wisconsin – Madison. PAL research is supported by NSF (IIS-2410981), the Wisconsin Alumni Research Foundation, and Sandia National Labs through a University Partnership Award.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: This paper claims (1) to introduce a formal model of thinking in RL environments (see Section˜3.2), (2) to formally show conditions when so-called thought actions will be chosen by an RL algorithm and how they could be helpful depending on policy initialization (see Section˜4), (3) to validate that some pre-trained LLMs satisfy the conditions when thought actions would be helpful (see Section˜5) and (4) to introduce a toy domain where multi-task pre-training leads to model-free RL discovering thinking as a useful strategy (see Section˜6).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Section˜7 discusses limitations as a starting point for future work. We also discuss a number of limitations throughout the paper, particularly when we introduce the thought MDP model in Section˜3.2.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All theoretical results are found in Section˜4 along with their proof. Key assumptions are stated along with the result or made in Section˜3.2.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Section˜5 completely specifies the LLM experiments and example prompts are provided in Appendix˜D. Section˜6 describes the toy-domain experiments and additional empirical details are found in Appendix˜E.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have included the code for the submission in the supplemental material. No external data sources are used.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: These details can be found in Section˜5, Section˜6, or Appendix˜E.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Confidence intervals are reported for all empirical results presented.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See Section˜6.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have read the Code of Ethics and confirm the paper conforms to them.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This paper is theoretical in nature and consequently has no immediate societal impact. Due to the dual-use nature of AI, by seeking to advance the field, our work will potentially contribute both to positive and negative societal impacts.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This is a theory paper with experiments on toy domains and does not release any new models that require safeguards.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the LLMs we use and also cite prominent software packages used (specifically Pytorch and Numpy).
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [No]
  64. Justification: No new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No human subjects or crowdsourcing used.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No human subject use.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: See Section˜5 and Appendix˜D.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.